OpenStudy (anonymous):

Ginny's class ordered 4 pizzas. They ate seven-twelfths of each pizza. How many pizzas did they eat? Answer 2 and one-third 3 and one-fourth 2 and one-fourth 1 and three-fourths

OpenStudy (anonymous):

2 and one third....\[4 \times \left(\begin{matrix}7 \\ 12\end{matrix}\right)\]